### v1

# Breaking the Ceiling of the LLM Community by Treating Token Generation as a Classification for Ensembling

Abstract
Ensembling multiple models has always been an effective approach to push the limits of existing performance and is widely used in classification tasks by simply averaging the classification probability vectors from multiple classifiers to achieve better accuracy. However, in the thriving open-source Large Language Model (LLM) community, ensembling methods are rare and typically limited to ensembling the full-text outputs of LLMs, such as selecting the best output using a ranker, which leads to underutilization of token-level probability information. In this paper, we treat the Generation of each token by LLMs as a Classification (GaC) for ensembling. This approach fully exploits the probability information at each generation step and better prevents LLMs from producing early incorrect tokens that lead to snowballing errors. In experiments, we ensemble state-of-the-art LLMs on several benchmarks, including exams, mathematics and reasoning, and observe that our method breaks the existing community performance ceiling. Furthermore, we observed that most of the tokens in the answer are simple and do not affect the correctness of the final answer. Therefore, we also experimented with ensembling only key tokens, and the results showed better performance with lower latency across benchmarks. [^1] Our code: https://github.com/yaoching0/GaC
Breaking the Ceiling of the LLM Community by Treating Token Generation as a Classification for Ensembling
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in a wide range of natural language processing tasks (Achiam et al., 2023; Touvron et al., 2023). Over time, new and more powerful LLMs are continually being released, pushing the boundaries of the LLM community (Meta, 2024; Alibaba, 2024). Due to the diversity of data sources, architectures and training methods, different LLMs have strengths and weaknesses in different tasks and contexts (Jiang et al., 2023). In addition to investing significant resources in training a superior LLM, ensembling multiple existing models is another effective way to break through the community performance ceiling (Huang et al., 2016), especially given the current trend in the open source LLM community to contribute only model weights rather than training data and procedures (AllenAI, 2024).
Figure 1: Motivation of GaC. The upper part shows CV classification ensemble, while the lower part illustrates ensemble at one text generation step.
Taking computer vision (CV) classification as an example, it is common to ensemble the output probability vectors of multiple models (e.g. by averaging) to achieve superior results (Krizhevsky et al., 2017). This approach remains effective even with recent CV models. As shown in Tab.1, we selected several common CV models (Chen et al., 2022; Tan and Le, 2019; Wang et al., 2022) for ensembling and observed better accuracy on ImageNet (Deng et al., 2009) compared to using a single model. Similarly, the popular decoder-only LLM architecture generates text by producing tokens one by one, with each generation step resulting in a probability vector of the length of the vocabulary. Inspired by CV, we propose to treat each generation step as a classification task, and by ensembling multiple models, we can achieve higher accuracy, as shown in Fig.1. There is already work that simplifies problems into binary tasks, exploiting the collective wisdom of LLMs and achieving better results (Schoenegger et al., 2024), demonstrating the feasibility of this approach.
CV Models on ImageNet Acc [%] ECE
EfficientNet-B1 78.55 0.072
RepGhostNet 78.81 0.053
PVTv2-B1 78.71 0.119
Ensembled CV Models (Averaged)
EfficientNet-B1 + PVTv2-B1 80.20$\uparrow$1.49 -
EfficientNet-B1 + RepGhostNet 80.06$\uparrow$1.25 -
EfficientNet-B1 + RepGhostNet
+ PVTv2-B1 80.62$\uparrow$1.81 -
LLMs on MMLU Acc [%] ECE
Llama-3-70b-Instruct 79.68 0.095
Qwen1.5-72b-chat 77.79 0.089
Yi-34B-Chat 72.75 0.090
Table 1: Performance of various CV models on ImageNet and LLMs on MMLU. $\uparrow$ indicates improvement over a single model.
Another advantage is that early errors in LLMs often snowball into later errors (Zhang et al., 2023). Ensembling during generation helps prevent the generation of inaccurate tokens at each step, thereby reducing misleading cues for subsequent token generation. In this paper, we conducted experiments at several points in time between November 2023 and June 2024, ensembling available state-of-the-art (SOTA) LLMs up to each of these points. We found that this approach significantly outperformed any single model available at those times on five popular benchmarks involving subject examination, mathematics, reasoning, and knowledge-based QA.
In addition, we found that for text generation it seemed unnecessary to ensemble every token. For example, for the question "What Andean animal has banana-shaped ears?" shown in Fig.1, the most critical part is for the LLM to generate the key token "llama". The initial part of the answer "It should be _" or "The animal is _" do not significantly affect the correctness of the final answer. Ideally, the step that produces the token "llama" is the one we want to ensemble.
Studies in CV classification have also shown that most samples are "simple" and can be correctly classified by most models (Wang et al., 2017), including cost-efficient ones, making the use of expensive models wasteful. To address this, CV classification used cascade inference Jazbec et al. (2024); Enomoro and Eda (2021), where a gate model passes a sample to a more powerful model only if its confidence falls below a threshold, thereby improving efficiency. Obviously, it is very important for cascading that the confidence of the gate model accurately reflects the accuracy. To ensure that LLMs are also suitable as gate models, we measured the Expected Calibration Error (ECE) (Guo et al., 2017) of CV models and LLMs on ImageNet and MMLU (Hendrycks et al., 2020), as shown in Tab.1. ECE is a metric that reflects the difference between a model’s confidence and its accuracy. We found that the ECE of CV models and LLMs were close. Therefore, in this paper, we also applied the cascade inference to LLMs by ensembling only the "key" tokens to speed up generation. Our experiments showed that this approach consistently achieved better performance with lower latency across different benchmarks.
2 Analysis and Prior Work
In this chapter, we review previous LLM ensemble studies and their features, as well as the problems our approach addresses. Previous studies can be categorized as follows:
Output-level ensemble methods select multiple candidate models and use their complete outputs for ensembling. Jiang et al. (2023) trained an additional ranking model (PairRanker) to score each candidate output and select the best one. Lu et al. (2023) and Shnitzer et al. (2023) trained a router to select the most appropriate candidate model given a question. However, these methods are limited to the existing candidate outputs and become ineffective if all the outputs are incorrect. Other studies have trained a fusion model to blend the outputs (Jiang et al., 2023; Wang et al., 2023b), overcoming the limitation of selecting only existing candidate outputs and often achieving superior results. However, the generalization of the fusion model is a major challenge, and they cannot fully exploit the probability information from each generation step.
Weight-level ensemble methods merge the weights of multiple models and are primarily used in multi-task learning (Yadav et al., 2024). The expectation is that the merged model will inherit capabilities across multiple tasks. However, a limitation is that the architectures of the models to be merged must be homologous, which limits the use of the capabilities of the LLM community. And it is rare to observe that the merged model outperforms the original models (Yu et al., 2023).
Training-level ensemble like FuseLLM (Wan et al., 2024) uses the output probability vectors of multiple models during training to ensemble as labels, rather than one-hot labels. In effect, this is a specific form of distillation that allows the model being trained to gain more information from the probability outputs of the ensembled (teacher) models. However, distillation is mainly used to improve small models, making it difficult to further improve the SOTA LLMs.
Our work can overcome the above limitations by ensembling at each generation step, allowing the output not to be confined to the original candidate output space and homologous architectures, while fully exploiting the probability information at each step. Our experiments in Sec.4 will also show that the ensemble consistently outperforms any single model, even SOTA LLMs. The main challenge, however, is that different LLMs typically have inconsistent vocabularies, leading to different dimensions in the probability vectors produced by different models. The most intuitive solution is to take the union of the vocabularies of the ensembled LLMs, denoted $V^{U}$, which includes all tokens from the participating models. Then, at each generation step, the output is first mapped to this union space $\mathbb{R}^{|V^{U}|}$ before ensembling.
Figure 2: The rate of identical tokenization for Oxford 5000 common words between different LLMs.
A potential problem with this approach is that different models may tokenize the same word differently, leading to conflicts. However, most mainstream LLMs use BPE or BBPE (Sennrich et al., 2015; Wang et al., 2020) to train tokenizers on sampled corpora, which tend to have similar sources (e.g. CommonCrawl) and distributions. This results in consistent tokenization for common words. For example, both Qwen1.5 and Llama3 (Bai et al., 2023; Meta, 2024) tokenize the word " alphabetically" into ["Ġalphabet", "ically"]. If both models intend to output this word, they will assign a higher probability to "Ġalphabet" first. We selected several popular LLMs (Young et al., 2024; Databricks, 2024; Almazrouei et al., 2023) and tokenized 5,000 commonly used English words (Oxford, 2018), and then calculated the proportion of identical tokenization results between each pair of LLMs, as shown in Fig.2. The proportion is above $90\%$ for all pairs, indicating that such conflicts can be ignored in most cases.
3 Proposed Method
In this section, we will first introduce the overall ensemble process of our GaC framework, and then explain the details in the following subsections.
Figure 3: Overview of GaC. The left side shows the creation of the mapping matrix, and the right side shows the ensembling during text generation with two LLMs.
3.1 Overall Process of GaC
When generating text, LLMs output a probability vector of the same dimension as their vocabulary. Given $n$ LLMs to be ensembled, we first take the union of their vocabularies and create a mapping matrix that can project the probability vectors to the union dimensions (Sec.3.2). At each generation step, all LLMs produce outputs that are mapped to the union vocabulary dimensions and ensembled to sample the next token. The tokenizer of each LLM then converts the sampled token into token IDs for the next step (Sec.3.3). As mentioned in Sec.1, not all tokens have the necessity for ensembling, so we also try to ensemble only certain key tokens (Sec.3.4).
3.2 Creating the Union Mapping
Given $\{\textit{LLM}_{1},\textit{LLM}_{2},\ldots,\textit{LLM}_{n}\}$ to ensemble, with their respective vocabularies $\{V^{1},V^{2},\ldots,\allowbreak V^{n}\}$, we first take the union of the vocabularies:
$V^{U}=\bigcup_{i=1}^{n}V^{i}.$ (1)
During this process, we record the positions of tokens from $V^{i}$ in $V^{U}$ and create corresponding mapping matrices $\mathbf{M}^{i}\in\mathbb{R}^{|V^{i}|\times|V^{U}|}$.
3.3 GaC Ensembling
At the start of text generation, we convert the input prompt into token ID sequences for each LLM. We denote the tokenizer of $\textit{LLM}_{i}$ as $\mathcal{T}^{\,i}:\textit{text}\rightarrow(\tau_{1},\tau_{2},\ldots,\tau_{m})$, which converts the input text into a sequence of token IDs. We calculate:
$\mathcal{I}^{\,i}=\mathcal{T}^{\,i}(\textit{prompt})\quad\text{for }i=1,\ldots,n$ (2)
where $\mathcal{I}^{\,i}$ is the input token ID sequence for $\textit{LLM}_{i}$.
For each generation step, we input $\mathcal{I}^{\,i}$ into $\textit{LLM}_{i}$ to obtain $p^{i}(\cdot\,|\,\mathcal{I}^{\,i})\in\mathbb{R}^{|V^{i}|}$, which represents the probability vector for the next token. These vectors are then mapped to the union vocabulary dimensions and averaged:
$q(\cdot)=\frac{1}{n}\sum_{i=1}^{n}p^{i}(\cdot\,|\,\mathcal{I}^{\,i})\cdot\mathbf{M}^{i},$ (3)
where $q(\cdot)$ is the ensemble probability vector. In Sec.4.3, we experimented with different ensemble weights and decided to use the average. We then sample a token $x\sim q(\cdot)$ as the result of this step. Finally, the sampled token is converted back into token IDs for each LLM and appended to $\mathcal{I}^{\,i}$:
$\mathcal{I}^{\,i}\leftarrow\mathcal{I}^{\,i}{}^{\frown}\mathcal{T}^{\,i}(x)\quad\text{for }i=1,\ldots,n$ (4)
We repeat (3) and (4) until the stopping criteria are met, such as outputting an end-of-sentence token or reaching the maximum length, as shown in Fig.3. In our implementation, different LLMs run in parallel on different GPUs, so the duration of each step is equal to the time taken by the slowest LLM. Since we have not modified a complete forward pass, our approach is compatible with techniques such as vLLM, DeepSpeed, quantization, and hardware optimizations (Kwon et al., 2023; Rasley et al., 2020).
3.4 Ensembling Key Tokens with Threshold
As mentioned in the last part of Sec.1, most tokens do not significantly affect the correctness of the response. From Tab.1, we can see that LLMs and CV models have similar ECE levels, suggesting that the confidence scores of LLMs may reflect accuracy to some extent. Therefore, we also experiment with ensembling only the tokens with confidence below a threshold $t$. We choose a model as the gate, denoted $\textit{LLM}_{g}$, and use its maximum probability at each step as the confidence score. During the ensemble, we replace the original (3) with:
$q(\cdot)=\begin{cases}\frac{1}{n}\sum_{i}p^{i}(\cdot\,|\,\mathcal{I}^{\,i})\cdot\mathbf{M}^{i}&\textit{if }\max(p^{g}(\cdot\,|\,\mathcal{I}^{\,g}))\leq t\\ p^{g}(\cdot\,|\,\mathcal{I}^{\,g})\cdot\mathbf{M}^{g}&\textit{otherwise}.\end{cases}$
Note that apart from $\textit{LLM}_{g}$, the other LLMs are not computed at every step, so their KV caches become stale. While there has been research using partial KV caches (Barad et al., 2023), for simplicity our work disables the KV caches of all LLMs except $\textit{LLM}_{g}$. This is an area for improvement and is listed in our future work.
4 Experiments
4.1 Overview
In this section, we first present the experimental setup, including the benchmarks and hardware used (Sec.4.2). We then test the effects of different ensemble weights for GaC (Sec.4.3) and compare it with other methods (Sec.4.4). We also select SOTA LLMs available at different times for ensembling to explore the performance ceiling at each time period (Sec.4.5). Finally, we experiment with thresholded ensembling to explore variations in latency and performance (Sec.4.6).
4.2 Experimental Settings
Benchmarks. GaC is not limited to specific tasks, so we tested it as broadly as possible. We selected a total of five benchmarks. For general capabilities, we chose MMLU (Hendrycks et al., 2020). For maths, we utilized GSM8K (Cobbe et al., 2021). For reasoning, we employed BBH (Suzgun et al., 2023). For knowledge capabilities, we included TriviaQA (Joshi et al., 2017) and NaturalQuestions (NQ) (Kwiatkowski et al., 2019). Note that all scores, including those for individual models, were computed locally under the same environment to ensure fairness [^2] Please see Appendix.A for more benchmarks details., using lm-evaluation-harness [^3] https://github.com/EleutherAI/lm-evaluation-harness v0.4.1 Gao et al. (2023).
Hardware and Latency. Each LLM was loaded on 1(n) A100 GPU(s) according to its memory requirements [^4] We listed each model and its hardware in Appendix.B., using naive model parallelism (HuggingFace, 2024) without optimization for inference. During ensembling, different LLMs were loaded on separate GPU(s) and executed in parallel, managed and communicated via Ray (Moritz et al., 2018). We also recorded the latency (ms/token). Each model performs a "dry run" after being loaded onto the GPU, generating 1024 tokens to warm up CUDA before experimentation, following the practice of Mehta et al. (2024).
4.3 Different Ensemble Weights
Model Acc [%] ECE
OpenChat-3.5-0106 64.53 0.0833
Qwen1.5-14B-Chat 67.20 0.1312
SOLAR-10.7B-Instruct-v1.0 64.48 0.2884
Yi-34B-Chat 72.75 0.0903
Qwen1.5-32B-Chat 75.12 0.1003
Nous-Hermes-2-Mixtral-8x7B-DPO 72.65 0.0789
Table 2: Two sets of LLMs of different sizes on MMLU. Smaller models on top, larger models on bottom.
Figure 4: Results of GaC ensemble with different weights for models from Tab.2. Smaller models ensembles on top, larger ones on bottom. The x-axis shows names participating in the ensemble (abbreviated).
Before proceeding with further experiments, we tested different ensemble weights for GaC. We used a simple averaging of the probabilities from each model in Eq.3. We now replace Eq.3 by $\frac{1}{\sum_{i}w^{i}}\sum_{i}w^{i}p^{i}(\cdot\,|\,\mathcal{I}^{\,i})\mathbf{M}^{i}$, where $w^{i}$ is the ensemble weight for $\textit{LLM}_{i}$. We set $w^{i}$ separately to each LLM’s score on MMLU, the MMLU score minus the ECE and 1 (i.e. averaging). We selected two sets of LLMs (Wang et al., 2023a; Kim et al., 2023; Young et al., 2024; Bai et al., 2023; NousResearch, 2024a) listed in Tab.2 with different sizes for GaC ensemble. The results on MMLU are shown in Fig.4. We observed no significant differences between the different weights, so for simplicity we decided to use averaging.
4.4 Comparison with Other Methods
Id Models MMLU GSM8K BBH TriviaQA NQ Avg. Latency
1 openchat_3.5 63.87 68.46 47.96 68.12 29.75 55.63 28.01ms/token
2 Nous-Hermes-2-SOLAR-10.7B 64.88 72.86 49.92 71.33 32.21 58.24 50.89ms/token
Ensemble Results for the Above Two Models (openchat and SOLAR)
3 LLM Blender (PairRanker) 64.23 74.07 50.18 70.55 32.55 58.32 57.86ms/token
4 OAssistRM 64.87 72.93 49.02 70.06 31.64 57.70 55.62ms/token
5 UltraRM 64.94 75.51 50.65 71.23 32.03 58.87 74.51ms/token
6 GaC (ours) 66.51 74.30 51.19 72.50 33.82 59.66 51.32ms/token
7 FuseLLM 63.94 65.50 46.32 64.57 29.06 53.88 28.31ms/token
[…]
9 Mixtral-8x7B-Instruct-v0.1 70.89 66.82 49.84 76.54 34.35 59.69 96.64ms/token
10 Yi-34B-Chat 72.75 68.76 50.88 70.01 29.81 58.44 67.96ms/token
Ensemble Results for the Above Two Models (Mixtral and Yi)
11 LLM Blender (PairRanker) 72.69 69.59 51.70 72.37 32.24 59.72 105.21ms/token
12 OAssistRM 73.34 70.15 51.91 72.79 30.69 59.78 99.75ms/token
13 UltraRM 69.49 71.09 52.27 73.82 32.36 59.81 114.57ms/token
14 GaC (ours) 74.83 71.21 52.64 75.60 33.52 61.56 98.13ms/token
Table 3: Results of comparison with other methods. Upper and lower halves represent different ensemble combinations. Blue indicates the best result for each ensemble. For id 8, GaC bottom right shows ensembled token proportion, top right shows threshold.
Baselines. We compared GaC with existing methods. First, we considered LLM Blender (Jiang et al., 2023), which employs PairRanker to rank the outputs of candidate LLMs and GenFuser to fuse these outputs. However, we found that GenFuser refused to answer a significant proportion of the questions in our chosen benchmarks. We therefore only used PairRanker to ensure fairness. We also included other rankers, such as OAssistRM [^5] https://huggingface.co/OpenAssistant/reward-model (Köpf et al., 2024) and UltraRM (Cui et al., 2023), which we ran in parallel on the GPUs hosting the ensemble LLMs to ensure low latency. These rankers scored the outputs and selected the best answers. Furthermore, we included the FuseLLM (Wan et al., 2024) (OpenChat-3.5-7B-Solar [^6] https://huggingface.co/FuseAI/OpenChat-3.5-7B-Solar), which uses probability information from multiple models during training for distillation.
Models for Ensemble. We chose two sets of models of different sizes. The smaller models included openchat-3.5 (Wang et al., 2023a) and Nous-Hermes-2-SOLAR-10.7B (NousResearch, 2024b) (teacher models for OpenChat-3.5-7B-Solar distillation). Larger models included Mixtral-8x7B-Instruct-v0.1 and Yi-34B-Chat (Jiang et al., 2024; Young et al., 2024).
Experimental Results. We ensemble the above two sets of LLMs using both our and baseline methods, and present the results in Tab.3. Our method showed superior performance with the lowest latency for both combinations (row ids 6 and 14). For row id 8, we used openchat-3.5 as the gate model with a threshold of 0.5 (Eq.5), resulting in only $7.68\%$ of tokens being ensembled. This slightly increased the latency from 28.01 to 31.34 ms/token, but achieved a performance close to that of SOLAR-10.7B (average score of 58.01 vs. 58.24), whose latency is 50.89 ms/token, further demonstrating the effectiveness of our method.
4.5 Breaking the Ceiling
Id Models MMLU GSM8K BBH TriviaQA NQ Avg. Date Latency
1 Yi-34B-Chat 72.75 68.76 50.88 70.01 29.81 58.44 2023/11/08 67.96ms/token
2 Mixtral-8x7B-Instruct-v0.1 70.89 66.82 49.84 76.54 34.35 59.69 2023/12/11 96.64ms/token
3 Qwen1.5-72B-Chat 77.79 83.33 48.94 65.69 27.02 60.55 2024/02/04 102.11ms/token
4 Llama-3-70B-Instruct 79.68 90.00 57.13 79.12 35.57 68.30 2024/04/18 150.32ms/token
5 Qwen2-72B-Instruct 82.30 89.70 62.57 73.58 33.11 68.25 2024/06/07 113.91ms/token
Ensemble the Above Models with GaC
6 Yi + Mixtral 74.83 71.21 52.64 75.60 33.52 61.56$\uparrow$3.13~2023/12/11 98.13ms/token
7 Qwen1.5-72B + Yi 79.83 77.27 52.05 70.88 33.80 62.77$\uparrow$3.65~2024/02/04 103.69ms/token
8 Qwen1.5-72B + Mixtral 79.55 75.76 54.19 75.71 31.09 63.26$\uparrow$4.47~2024/02/04 112.83ms/token
9 Llama-3 + Qwen1.5-72B 81.49 87.06 56.73 78.60 36.01 67.98$\downarrow$0.47~2024/04/18 153.96ms/token
10 Qwen2-72B + Llama-3 83.54 90.91 63.99 79.29 37.65 71.08$\uparrow$4.06~2024/06/07 151.56ms/token
Task-specific Top-2 Model Ensemble with GaC
11 Top-2 (~2024/04/18) 81.49 87.96 58.64 80.84 37.95 69.38$\uparrow$1.58~2024/04/18 -
12 Top-2 (~2024/06/07) 83.54 90.91 63.99 80.84 37.95 71.45$\uparrow$4.61~2024/06/07 -
Table 4: Ensemble of available SOTA LLMs from different periods. The top part lists the individual models, while the bottom part shows the ensemble results (model names abbreviated). $\uparrow$ indicates the percentage improvement over the individual models.
Figure 5: Ensemble results of CV models with different accuracy gaps on ImageNet. Models’ accuracies are next to their names. Each cell shows the ensemble accuracy, with the improvement over the best single model in parentheses.
In this experiment, we aimed to break the performance ceiling of the open source LLM community at different times. We chose the SOTA LLMs released between November 2023 and June 2024, as listed in the upper part of Tab.4, excluding models with more than 100 billion parameters due to hardware limitations. We then ensemble the SOTA LLMs available at different times, as shown in row ids 6-10, and observe an improvement of 3.13% to 4.47% over the best single model at each time. An exception is 2024/04/18, when Llama-3-70B-Instruct was released and significantly improved performance over the previous SOTA LLMs (average score increased from 60.55 of Qwen1.5-72B-Chat to 68.30), resulting in a drop in performance after ensemble due to the large gap.
However, with the release of Qwen2-72B-Instruct, which showed comparable performance to Llama-3-70B-Instruct, the ensemble again led to significant improvements (row id 10). In rows 11 and 12, we ensemble the top two best-performing models for each benchmark at the two most recent times, including the challenging time of 2024/04/18, and observe performance gains with this task-specific top-two ensemble even on 04/18 (row id 11). Finally, row id 12 shows the best results available for the open source community on 2024/06/07. By pushing the boundaries of the community, we can narrow the gap with proprietary models and promote the democratization of LLMs.
Since ensembling models with large performance differences could lead to performance degradation (row id 9 in Tab.4), we also tested this hypothesis with CV models. We ensemble PVTv2-B1 Wang et al. (2022) with different sizes of EfficientNet (Tan and Le, 2019) on ImageNet (Deng et al., 2009) by averaging their outputs, as shown in Fig.5. We observed that as the accuracy gap between the two models increased, the ensemble gains decreased and eventually became negative. This suggests that it is advisable to ensemble models with similar levels of performance.
4.6 Ensemble with Threshold
Id Models Threshold MMLU GSM8K BBH TriviaQA NQ Avg. Latency
1 Llama-3-70B-Instruct - 79.68 90.00 57.13 79.12 35.57 68.30 150.32ms/token
2 Llama-3-8B-Instruct - 65.08 76.26 44.72 67.67 26.48 56.04 34.30ms/token
3 Phi-3-mini-4k-instruct - 67.11 79.00 47.26 55.78 17.98 53.43 31.82ms/token
4 Qwen1.5-72B-chat - 77.79 83.33 48.94 65.69 27.02 60.55 102.11ms/token
5 Qwen1.5-32B-Chat - 75.12 75.97 53.89 62.57 22.96 58.10 59.01ms/token
Ensemble with threshold to match Qwen1.5-72B-Chat performance (avg. 60.55)
6 Llama-3-8B + Llama-3-70B9.30% 0.5 69.51 82.86 47.65 74.35 33.74 61.62 68.97ms/token
7 Qwen1.5-32B + Qwen1.5-72B6.31% 0.5 75.53 81.82 55.87 63.91 27.69 60.96 77.86ms/token
Ensemble with threshold to match Qwen1.5-32B-Chat performance (avg. 58.10)
8 Llama-3-8B + Llama-3-70B6.98% 0.45 68.06 81.79 46.66 73.43 33.74 60.74 58.99ms/token
9 Phi-3 + Llama-3-70B7.59% 0.5 68.46 78.57 50.34 69.08 29.51 59.19 51.61ms/token
Table 5: Thresholded ensemble results. The top lists individual models, while the bottom shows ensemble combinations (model names abbreviated). The percentage in the bottom right of the combination names represents the proportion of tokens ensembled.
Models Threshold MT-Bench Latency
Llama-3-8b-Instruct - 8.03 33.75ms/token
Llama-3-70b-Instruct - 8.80 128.50ms/token
Qwen1.5-72b-chat - 8.33 87.39ms/token
GaC (Llama-3-8b + Llama-3-70b)10.10% 0.55 8.34 69.28ms/token
Qwen1.5-32b-chat - 8.12 58.49ms/token
GaC (Llama-3-8b + Llama-3-70b)7.24% 0.5 8.16 57.28ms/token
Table 6: Thresholded ensemble on MT-Bench. GaC shows the ensemble combinations (model names abbreviated), with the proportion of tokens ensembled shown at the bottom right.
In this experiment, we used the thresholded ensemble (Sec.3.4) to explore variations in latency and performance. We selected models of different sizes (Abdin et al., 2024; Meta, 2024; Bai et al., 2023), listed in the upper part of Tab.5, pairing a smaller model with a larger model and using the smaller model as the gate model for the ensemble. We aimed to match the performance (average score) of Qwen1.5-72B-chat and Qwen1.5-32B-chat with our ensemble, but with lower latency, and the results are shown in row ids 6-9. Interestingly, even when combining the two Qwen models themselves with a threshold of 0.5 (row id 7), where 6.31% of the tokens were ensembled, we observed slightly higher performance than Qwen1.5-72B-chat (average score increased from 60.55 to 60.96) and lower latency (102.11 to 77.86 ms/token). We believe this is a promising new way to speed up inference.
We also observed lower latency in row ids 6, 8 and 9 of Tab.5 with comparable performance to Qwen1.5-72B-chat or Qwen1.5-32B-chat. In addition, similar trends were observed in the MT-Bench (Zheng et al., 2023) using LLM as judge (GPT-4-0613) in Tab.6, with scores calculated using FastChat [^7] https://github.com/lm-sys/FastChat in our local environment. For the combinations of Llama-3-8B-Instruct or Phi-3-mini-4k-instruct with Llama-3-70B-Instruct, we directly adopted the output of the larger model if the probability of the smaller model was below the threshold. This is based on our observations in Fig.5, where ensembling models with large performance gaps resulted in reduced performance.
5 Conclusion
In this paper, we present a token-level ensembling framework called GaC, which fully exploits the probability information at each generation step. In our experiments, we have surpassed the performance ceiling of open-source SOTA LLMs available at different time periods (Sec.4.5), further narrowing the gap between open-source and proprietary models. This progress promotes the democratization of LLMs and provides new motivations for future research, enabling better exploitation of collective intelligence. In addition, we experimented with ensembling just a few tokens and found that this approach can achieve better performance with lower latency (Sec.4.6), opening up new avenues for accelerating inference.
Contemporaneous Works
We have noticed several contemporaneous works related to our research, all of which aim to address the vocabulary discrepancy between different models. Xu et al. (2024) proposed EVA, which trains a projection matrix between each pair of LLMs, using the overlapping tokens from their vocabularies as a bridge. DEEPEN (Huang et al., 2024) converts the output probabilities to a relative representation using anchor tokens before ensembling, and then inverts back to the original model’s vocabulary space using gradient descent, which requires an additional $7\%$ to $29\%$ of time per generation step. In contrast, our method requires no additional training and only a single matrix multiplication and tokenization for each model during ensembling, with minimal time cost.
Limitation
Like other ensemble methods, the approach proposed in this paper requires more computational resources. Although different models can be run in parallel on separate GPUs, so that latency only depends on the slowest model, the overall computational load is additive, raising the threshold for use.
Appendix A Benchmarks
This paper uses the lm-evaluation-harness v0.4.1. The task names in the repo corresponding to each of the benchmarks we used are as follows:
MMLU: mmlu_flan_n_shot_generative, 5-shots
GSM8K: gsm8k, 5-shots.
BBH: bbh_fewshot, 3-shots.
TriviaQA: triviaqa, 5-shots.
NQ: nq_open, 5-shots.
Appendix B Hardware Specifications
Models Hardware
Phi-3-mini-4k-instruct 1x A100
Llama-3-8B-Instruct 1x A100
openchat_3.5 1x A100
OpenChat-3.5-0106 1x A100
Qwen1.5-14B-Chat 1x A100
SOLAR-10.7B-Instruct-v1.0 1x A100
Nous-Hermes-2-SOLAR-10.7B 1x A100
Yi-34B-Chat 1x A100
Qwen1.5-32B-Chat 1x A100
Nous-Hermes-2-Mixtral-8x7B-DPO 2x A100
Mixtral-8x7B-Instruct-v0.1 2x A100
Llama-3-70B-Instruct 2x A100
Qwen1.5-72B-Chat 3x A100
Qwen2-72B-Instruct 3x A100
Table 7: Models and Hardware